OpenStudy (anonymous):

Nicotine contains 74.0% C and 8.7% H and 17.3% N and has a mass of 162.1 grams What's the molecular formula?

OpenStudy (welshfella):

first find the ratio of the atoms in the molecule by dividing each percentage by the ralative atomic mass of the element eg C = 74 / 12

OpenStudy (welshfella):

H = 8.7 / 1.008 N = 17.3 / 14.007

OpenStudy (welshfella):

divide each of the results by the smallest to give values very close to whole numbers You will then have the ratio of the elements in the whole numbers.

OpenStudy (welshfella):

THis is the empirical formula for nicotine Add up the atomic masses to get the molecular mass Compare this with the ginen mass 162.1 From this you get the molecullar formula.